35 ml sample of hydrogen peroxide gives off 500ml of o2 at 27°C and 1atm pressure. volume strength of h2o2 sample wii be

Grade:11

2 Answers

Arun
6 years ago
Dear Avinash
 
Use PV=nRT and solve for n = PV/RT to get the number of moles of O2 given off 
and substitute 
n=PV/RT 
Substitute 
P= 1 atm 
V=.500 L (notice you need to convert to L to match the units in R) 
R=0.082 L atm/ (mol K) 
T= 300 K 

a. Solve for n to get moles of O2 released. n=.0203 

b. You can get the volume of O2 relesed at stp from 22.4 L/moles of O2 at STP = 0.4553 

c. For every mole of H2O2 you get one mole of O2 released, so that tells you 

moles H2O2 in the solution = 00203 = 15.48 g H2O2 in the solution 

If you assume the density of H2O2 is the same as water, then you have 15.5 mL of H2O2 in the solution and the V/V % is 15.5x100/35 = 44% You could also say the w/v % is the same number. It is unclear which you are looking for.
 
 
Regards
Arun (askIITians forum expert)
Dylan
5 years ago
See, first find volume if oxygen that canbe produced at stp...
Therfore, v1/t1=v2/t2(modi 's law)
V2=500*273/300(by substituting values in the equation)
=455ml
35ml can produce 455ml of oxygen at stp,
Therfore(unitary method) 1000ml can produce 1000*455/35
=13000ml=13l